OpenStudy (dbour):

FAN+MEDAL Given acute triangle ABC, find the area with the given information: Altitude (height): 7.5

OpenStudy (wolf1728):

Do you have more information?

OpenStudy (dbour):

@wolf1728 Unfortunately not, otherwise I may have been able to solve this myself. I've checked the question/steps numerous times, and no further information is provided beside the altitude/height. Triangle ABC is acute, with Segment BD (height) being perpendicular to segment AC.

OpenStudy (dbour):

@wolf1728 Thank you for taking interest, btw! :-)

OpenStudy (dbour):

It's possible that it's just written poorly, and it actually wants me to write an equation to solve it as if I was given more information/measurements/angles.

OpenStudy (alexandervonhumboldt2):



OpenStudy (alexandervonhumboldt2):

this is all what is given? starnge

OpenStudy (wolf1728):

The best I could say is triangle area = (1/2) * base * height area = (1/2) *base * 7.5 area = 3.75 * base

OpenStudy (dbour):

@AlexandervonHumboldt2 I completely agree! Very odd...

OpenStudy (alexandervonhumboldt2):

maybe we can use law of sines or cosines let me see

OpenStudy (alexandervonhumboldt2):



OpenStudy (dbour):

@wolf1728 That makes sense given the available information. However, this is in the unit of trigonometric ratios, so I believe it expects me to use the law of sines / cosines as @AlexandervonHumboldt2 just pointed out :)

OpenStudy (wolf1728):

There are 2 other triangle area formulas but they require knowing all 3 sides or 2 sides and the included angle

OpenStudy (wolf1728):

Area = (1/2) * side 1 * side 2 * sine (A) where A is the included angle

OpenStudy (dbour):

@wolf1728 I think you're on to something. It says I should be able to use the formula A = 1/2 a X b X sinC

OpenStudy (alexandervonhumboldt2):

law of sines don't seem to be usable let me see if cosines.......

OpenStudy (alexandervonhumboldt2):

there is not enough info for using both sines and cosines there should be another method

OpenStudy (wolf1728):

As I said there are 3 formulas for finding triangle area but we don't have enough information to use any of them.

OpenStudy (dbour):

The frustrating part is that the tool the problem is given on can give me all the angle measures / side lengths, but that would be skipping the steps.

OpenStudy (wolf1728):

If you know the angles that would provide a solution.

OpenStudy (dbour):

@wolf1728 I agree. I'm just going to write the solution to finding the area with the given information without solving it, as it is literally impossible to solve without more information. Thank you both @wolf1728 @AlexandervonHumboldt2 for your wonderful help! Have a wonderful night you guys!

OpenStudy (wolf1728):

u r welcome - glad to help

OpenStudy (alexandervonhumboldt2):

:) however this problem will confuse me all night lol